OpenStudy (anonymous):

x^2+xy-y^2=4 implicit differentiation?

OpenStudy (dan815):

diff wrt to wat?

OpenStudy (anonymous):

Find by implicit differentiation

OpenStudy (dan815):

lemme show u something when your nomally have \[y=f(x)\\ and~ you~ do~ \frac {d}{dx} ~to ~both ~sides\\ \frac {d}{dx}(Y)=\frac {dy}{dx}=\frac {df}{dx}\\ similarily\\ \frac {d}{dx}(f(y))=\frac {df(y)}{dx} * \frac {dy}{dx}, ~ by~the ~chain ~rule\]

OpenStudy (dan815):

follow these rules and differentiate

OpenStudy (dan815):

suppose y=g(x)

OpenStudy (dan815):

rewrite your whole equation like this

OpenStudy (dan815):

x^2xy-y^2=4 x^2x*g(x)-g(x)^2=4 now let side is just some function of x, apply chainrule and other rules you know to differentiate wrt to x

OpenStudy (anonymous):

no i had a typo my bad

OpenStudy (anonymous):

i was supposed to have a +* in between x^2 and xy

OpenStudy (dan815):

kk just carry on same rules apply

OpenStudy (dan815):

where ever you see a f(y) you have to do f'(y) * y'

OpenStudy (dan815):

and then for the f(x) you are doing the samething as before

OpenStudy (dan815):

and for the right side the any constant differentiated wrt to any variable is just zero

OpenStudy (anonymous):

so 2x+xy'-2y*y'=0

OpenStudy (dan815):

product rule for xy

OpenStudy (dan815):

d/dx (xy)=1*y + x*y'

OpenStudy (dan815):

because d/dx ( x*f(x)) = 1*f(x) + x*f'(x)

OpenStudy (anonymous):

so when ever you have ;like lets say 4xy we use the product rule

OpenStudy (dan815):

yah you think of 4x as a function and y as af unction

OpenStudy (dan815):

d/dx (f(x) G(x)) = f'*(x) *G(x) +f(x) * G'(x)

OpenStudy (anonymous):

ok i get it now i got y'=2x+y/2y-x

OpenStudy (dan815):

ok

OpenStudy (anonymous):

Refer to the three statement solution using the Total Derivative with Mathematica 9.

OpenStudy (perl):

@robtobey mathematica = wolfram?

OpenStudy (anonymous):

Yes. Special pricing for students and I believe faculty. Current version is Mathematica 10. http://www.wolfram.com/mathematica/pricing/students.php